OpenStudy (anonymous):

HELP ME PLEASE 2^3x-1 = 4^x

OpenStudy (anonymous):

Do you want to work from the same textbook?

OpenStudy (anonymous):

\[2^{3x-1}= 4^{x}\]

OpenStudy (anonymous):

Just curious because I'm not positive, but doesn't 4^x need to get changed to 2(2^x)?

jimthompson5910 (jim_thompson5910):

close DonaldRoyMiller, but not quite

jimthompson5910 (jim_thompson5910):

since 4 = 2^2 we can say this.... \[\large 2^{3x-1}= 4^{x}\] \[\large 2^{3x-1}= (2^2)^{x}\] \[\large 2^{3x-1}= 2^{2x}\]

OpenStudy (anonymous):

They need a common base number, right. OH< HERE WE GO AGAIN WITH SOMEONE JUST GIVING AN ANSWER !!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

nope, I didn't give the answer since I didn't finish (you need to solve for x)

jimthompson5910 (jim_thompson5910):

so please don't jump to conclusions and actually read what I posted

OpenStudy (anonymous):

i SEE. Okay. You're right. But it's frustrating at how often that happens.

OpenStudy (anonymous):

how do i solve for x??

OpenStudy (anonymous):

What you do is ignore the base for now and focus on the exponent.

OpenStudy (anonymous):

You aren't going to download the free textbooks are you.

OpenStudy (anonymous):

solve 3x-1 = 2x

OpenStudy (anonymous):

do i add the 1 to the 2?

OpenStudy (anonymous):

You need to know the process, or you're going to have this same problem over and over again. Once you know how it works, you're good to go.

OpenStudy (anonymous):

3x and 2x are two numbers being multiplied. So you have to work with likes. The one and the two are constants. They don't change. Know what I mean?